OpenStudy (anonymous):

A transformer has 175 turnings on the primary coil and 3000 turnings on the secondary coil. If the output voltage is 4000 V, what is the input voltage?

OpenStudy (anonymous):

\[\frac{ V_p }{ V_s }=\frac{ N_p }{ N_s }\rightarrow V_p= \frac{ N_p }{ N_s }V_s=???\]